1. The number of slurms in Wormulon is 200 in year 1600, 400 in 1800, and 600 in 1820. (1) [2 points] Compute the percentage change in slurms over the two periods defined by these three dates. (2) |3 points] Compute the average rate of growth in slurms for the period 1600-1800, 1800-1820, and 1600-1820. 3) lB point] Your answer for the growth rate in the period 1600-1820 in question (2) is not equal to the growth rate in either sub-period. Comment
Answer #1

1. Percentage change in slurms over the period 1600-1800 = [(Slurms in 1800 - Slurms in 1600)/Slurms in 1600]*100
= [(400 - 200)/200]*100 = (200/200)*100 = 100%
Percentage change in slurms over the period 1800-1820 = [(Slurms in 1820 - Slurms in 1800)/Slurms in 1800]*100
= [(600 - 400)/400]*100 = (200/400)*100 = 50%

2. Average rate of growth = Percentage change/Number of years
Average rate of growth for the period 1600-1800 = Percentage change over the period 1600-1800/Number of years = 100%/200 = 0.5%

Average rate of growth for the period 1800-1820 = Percentage change over the period 1800-1820/Number of years = 50%/20 = 2.5%

Average rate of growth for the period 1600-1820 = Percentage change over the period 1600-1820/Number of years
Percentage change over the period 1600-1820 = (Slurms in 1820 - Slurms in 1600)/Slurms in 1600]*100
= [(600 - 200)/200]*100 = (400/200)*100 = 200%
So, Average rate of growth for the period 1600-1820 = Percentage change over the period 1600-1820/Number of years = 200%/220 = 0.91%

3. Growth rate in the period 1600-1820 is not equal to the growth rate in either sub period because of difference in the number of years that the percentage growth is divided by to calculate the average growth rate.


​​​​